For each of the following situations, calculate the z-statistic (z) make a decision about the null hypothesis (reject, do not reject), and indicate the level of significance(p>.05, p<.05, p< .01) = 8.00: µ =16; σ = 4.00: z = -2, reject H0 ( 2 standard deviations), (p>.05) =39.54; µ =34.22; σ = 2.18: z = 2.44, reject H0( more than 2 standard deviations) (p>.05) = 1.19; µ = .92; σ =. 17: z = 1.59, reject H0 (1.5 standard deviations) ( p>.05) = 56.81; µ = 64.45; σ = 2.68: z = -2.85, reject H0( more than 2.5 standard deviations) ( p>.05)
Answer #1

Ans:

a)

z=(8-16)/4

z=-2

reject H0(p<0.05)

b)

z=(39.54-34.22)/2.18

z=2.44

reject H0(p<0.05)

c)

z=(1.19-0.92)/0.17

z=1.59

Fail to reject the null hypothesis.(p>0.05)

d)

z=(56.81-64.45)/2.68

z=-2.85

Reject H0(p<0.05)